OpenStudy (dumbsearch2):

Use the Triangle Inequality to determine whether it is possible to draw a triangle with sides of the given measures: 16, 17, 20 Explain me the fastest way to do a problem like this one :p Thank you. And don't give me the answer to this one, just tell me how to do this stuff :3

OpenStudy (dumbsearch2):

From what I read up on Wikipedia about the Triangle Inequality, this would be false. Just want to verify this.

OpenStudy (reemii):

i think it's possible. here it would be impossible: if lengths are 10, 1 and 1... see why?

OpenStudy (anonymous):

16, 17, 20 ---------------- Hypotenuse: 20 Leg: 16 leg: 17

OpenStudy (aravindg):

Well the triangle inequality tells for a triangle with sides a,b,c \[a+b \ge c\]

OpenStudy (dumbsearch2):

Yeah. Maybe. Wiki is confusing... Yes, I wouldn't see why it wouldn't work, but from my limited understanding of this property, this is what I got from it: if two sides don't add up to be greater then the other remaining one, then it's false... Though this really sounds wrong...

OpenStudy (anonymous):

The inequality says that any two sides must add up to greater than the third side, so if the two shorter sides are longer than the longest side you are all good.

OpenStudy (anonymous):

a^2 + b^2 = c^2 16^2 + 17^2 = 20^2 256 + 289 = 400 545 = 400 FALSE a^2 + b^2 > c^2 16^2 + 17^2 > 20^2 256 + 289 > 400 545 > 400 TRUE a^2 + b^2 < c^2 16^2 + 17^2 < 20^2 256 + 289 < 400 545 < 400 FALSE Therefore, yes you can form a triangle with the side lengths given

OpenStudy (dumbsearch2):

huh. Explain a little moar ? :3

OpenStudy (anonymous):

well: a^2 + b^2 = c^2 is if it is a right triangle... which is not in this case...cuz it was false...

OpenStudy (aravindg):

look at theorem 1 http://www.regentsprep.org/Regents/math/geometry/GP7/LTriIneq.htm

OpenStudy (dumbsearch2):

GOT IT! So I had the theorem backwards :p Thanks guys for clearing this one up. And now I don't know who to medal.. :[

OpenStudy (anonymous):

just give it to whom you think deserves it :)

OpenStudy (dumbsearch2):

omg this is a hard one.

OpenStudy (dumbsearch2):

I mean, you gave a lot of info, @some_someone... but @AravindG gave me a SUPER helpful link...

OpenStudy (anonymous):

i gave one to @@AravindG so do not worry :)

OpenStudy (dumbsearch2):

g00d :)

OpenStudy (aravindg):

I gave a medal to @mrbarry :)

OpenStudy (anonymous):

see, we are all happy now ^_^ lol